# Reinforcement Unlearning via Group Relative Policy Optimization

### v1

@@ 5.3 Ablation Studies
PURGE is robust against 8 out of 9 adversarial attacks in the forget set. Figure 4 shows the difference between the original performance on the forget set and the unlearned models according to PURGE (black) vs. SoTA in 9 adversarial scenarios that aim to recover the unlearned knowledge. A negative difference here is preferred since the forget quality is better. As supported by the average scores in Table 1, notice how PURGE consistently unlearns the forget set and does not adversarially leak this information, because it is sufficiently suppressed in its internal latent representation.
@@
Figure 3: Performance of PURGE across Qwen model sizes. The FLU metric (left) is normalized for interpretability, where lower values indicate better performance. Percentage scores (right) are shown, where higher values are preferable
@@
Figure 4: PURGE (black) works consistently on adversarial attacks (See D.1 for details) over the baseline in Forget Quality % ($\downarrow$). We report difference of each method with the baseline performance. Hence, negative differences are better (unlearning works).
@@
PURGE maintains the model utility regardless of model size. To investigate how well our unlearning method scales with model size, we apply Purge to three checkpoints from the Qwen‑2.5 family (0.5B, 1.5B and 3B parameters) and evaluate them with a subset of the benchmarks introduced in Section 5.1. The results are summarized in Figure 3. Across all three forgetting probes (FB, QA and AA), we observe a monotonic incline in performance as the parameter count increases. This trend indicates that larger models retain unwanted information more stubbornly and are therefore harder to unlearn, which is consistent with recent evidence (Carlini et al., 2022) that memorization in LLMs scales super‑linearly with size. In contrast, the plot on the right of Figure 3 shows that useful capabilities are largely preserved after PURGE, independent of model scale, which is in line with what we expect to happen based on Theorem 2. All five utility metrics remain stable, indicating that PURGE effectively removes targeted knowledge without harming general downstream performance, even with millions of additional parameters.
@@ 6 Conclusion
In this work, we presented PURGE, a novel framework for targeted LLM unlearning that bridges the gap between preference-based approaches and the recent successes of RLVR. By formulating unlearning as a verifiable task and utilizing GRPO, PURGE removes designated facts while rigorously preserving the model’s linguistic capability and downstream utility. Our theoretical analysis provides formal guarantees of information suppression and high-probability bounds showing that forgetting generalizes beyond the prompts seen during training, while protecting overall model quality. Comprehensive experiments on the RWKU benchmark show, that PURGE outperforms most SoTA methods in fluency, adversarial robustness – achieving +5.48% higher fluency scores and a 12.02% gain in resistance to adversarial attacks – while preserving more than 98% of the baseline performance on standard NLP tasks and using up to $\times$46 less tokens per target compared to SoTA approaches. In the future, we want to refine the definition of unlearning targets by extracting semantically related concepts directly from the model’s embedding space, enabling even more precise, model-aware unlearning. We also aim to explore and apply our method to batch unlearning, enabling more realistic and scalable applications of unlearning.
@@ A.3 Putting Theorem 1 in GRPO
Notice that this does not mean that GRPO, after unlearning, has a zero probability of emitting forbidden tokens. This means that at the limit, this probability tends to zero, indicating that GRPO is a suitable approach to do unlearning.
@@ A.4 Proof of Theorem 2
Since the KL‐penalty in the GRPO objective explicitly bounds the average per‐prompt KL divergence, we can pull the square‐root bound outside.
@@ A.5 Regret‐to‐Retraining Theorem
Gist of Theorem 3: If you run GRPO for $T$ iterations with a standard $1/\sqrt{t}$ learning‐rate schedule, then on average your task‐loss will be within $O(1/\sqrt{T})$ of the loss achieved by fully retraining from scratch on the retained data, meaning that after a modest number of updates, PURGE’s performance on the kept tasks is essentially as good as a complete retrain, at a tiny fraction of the cost. Notice that this theorem holds, however we cannot measure it empirically since we are not able to train an LLM from scratch. Therefore, we do not include it in the main content of the paper. We invite researchers to develop empirical demonstrations of this theorem in LLM unlearning and to critically evaluate whether retraining from scratch truly constitutes the gold standard for unlearning in LLMs.
@@
We cast GRPO as mirror‐descent on the space of token‐probability distributions, using the KL divergence as the Bregman distance. Concretely, each update solves a proximal optimization of the form
@@ Appendix B Additional Experiments
B.1 Extended Results on Unlearning Finetuning Dynamics
Detailed Reward Analysis
Figure 5 shows the complete reward trajectories for each unlearning-target training run. At each training step, we record the mean reward across all model completions. Initially, most curves begin at a low mean reward, indicating that the model does not yet adhere to our defined reward function – see Equation 13 – and then steadily rise, demonstrating that the model progressively learns to satisfy our custom reward criteria, hence unlearn the target.
Figure 5: Training reward trajectories for each unlearning target plotted against the global training step. Each of the 20 subplots shows the full evolution of the model’s ability to follow the reward function for a specific target over the course of its training run, highlighting differences in convergence speed and stability across experiments. Note that Theorem 1 is satisfied on all targets.
Detailed KL Divergence Analysis
Figure 6: KL divergence trajectories for each unlearning target plotted against the global training step. Each of the 20 subplots shows the evolution of the KL divergence throughout its respective training run, illustrating variations in divergence reduction dynamics and convergence behavior across different unlearning targets. Notice that Theorem 2 is satisfied for all targets.
B.2 Additional TOFU Benchmark Results
To further evaluate the generalization capabilities of PURGE, we extended our experiments to the TOFU benchmark. Only minor adjustments to the forget–corpus construction were required, confirming that PURGE does not rely on RWKU-specific assumptions and generalizes to alternative unlearning settings. We follow the OpenUnlearning (Dorna et al., 2025) framework and reproduce all baselines on Llama-3.2-1B-Instruct, using the evaluation protocol and reference implementations described in the GitHub repository provided by OpenUnlearning (specifically the results reported in open-unlearning/docs/repro.md).
Table 2 summarizes the results across three key metrics: Forget Quality, Forget Truth Ratio, and Model Utility. Forget Quality and Forget Truth Ratio closer to 1 indicates stronger forgetting, while higher Model Utility indicate better preservation of model capabilities. PURGE remains competitive with contemporary approaches (e.g., GradDiff, RMU, UNDIAL), despite operating under a conceptually different mechanism. Although not state-of-the-art on TOFU, PURGE exhibits stable and robust performance, highlighting future opportunities for optimization and deeper theoretical analysis. In summary, these additional results demonstrate that PURGE maintains strong performance on TOFU with minimal methodological adjustments, underscoring its generality and revealing promising avenues for future refinement.
@@ C.1 Baseline Details
User: You are an AI Assistant who is supposed to forget about Stephen King and provide answers without its knowledge as if you never knew about it. Don’t tell anyone that you unlearned anything.
@@
Building on the rejection templates from the TOFU (Maini et al., 2024) benchmark, we curate 100 negatively‑labelled prompts and fine‑tune by minimizing
@@ C.2 Hyperparameter Selection
We tuned the number of epochs and the learning rate ($\eta$) per training stage to balance stability (avoiding catastrophic drift from the base Phi-3 model) and sufficient adaptation to each objective. In Table 3 below we summarize the choices and their rationale.
@@
Overall, we fixed the epoch count to 3 to simplify scheduling and monitoring, and we scaled $\eta$ inversely with the expected gradient volatility and risk of catastrophic forgetting for each stage.
@@ D.1 RWKU Dataset Sets
The goal of the forget set is to meeasure the unlearning efficacy—i.e., whether the model ceases to recall or use the specific knowledge designated for removal. It is partitioned into three splits, that test different aspects of unlearning:
@@
More specifically the AA split is comprised of 9 adversarial attack types, we give a few examples:
@@
The MIA set is used to detect whether the model still betrays membership of target fragments in training data, serving as a privacy proxy. It is comprised out of the following splits:
@@
The utility set is used to quantify the side effects that unlearning has on broad capabilities beyond local neighborhood knowledge. We use well-known benchmarks that test different LLM capabilities: (1) General Ability (Gen): MMLU (5-shot accuracy); (2) Reasoning Ability (Rea): BBH with 27 subtasks, CoT with 3-shot prompts (Exact Match); (3) Truthfulness (Tru): TruthfulQA MC1 (6-shot accuracy); (4) Factuality (Fac): TriviaQA (6-shot F1); and (5) Fluency (Flu): AlpacaEval, weighted bi-/tri-gram entropy.
@@ D.2 Metrics
Measures longest-common-subsequence overlap between the model’s output and the reference. Forget Set: Lower is better ($\downarrow$), less overlap means the target knowledge is removed. Neighbor Set: Higher is better ($\uparrow$), we want the model to retain nearby knowledge.
@@ Appendix F PURGE Unlearning Qualitative Examples
In order to illustrate the practical impact of our PURGE unlearning procedure, we present a series of qualitative examples drawn from three well-known public figures. Each example contrasts the model’s original (“Base Model”) output with its corrected (“PURGE”) output, highlighting how targeted removal of spurious knowledge leads to more accurate and robust responses, showcasing the versatility of the method in real‐world scenarios.

### v2

@@ 5.3 Ablation Studies
PURGE is robust against 8 out of 9 adversarial attacks in the forget set. Figure 4 shows the difference between the original performance on the forget set and the unlearned models according to PURGE (black) vs. SoTA in 9 adversarial scenarios that aim to recover the unlearned knowledge. A negative difference here is preferred since the forget quality is better. As shown by the average scores in Table 1, PURGE consistently unlearns the forget set and does not adversarially leak this information. By minimizing the probability of generating characteristic tokens associated with a concept, PURGE induces broader semantic suppression. Because PURGE penalizes entire completions (i.e., full reasoning trajectories), the model learns to reduce not only explicit mentions but also the contextual semantics in which those mentions typically arise.
@@
Figure 3: Performance of PURGE across Qwen model sizes. The FLU metric (left) is normalized for interpretability, where lower values indicate better performance. Percentage scores (right) are shown, where higher values are preferable.
@@
Figure 4: PURGE (black) works consistently on adversarial attacks (See D.1 for details) over the baseline in Forget Quality % ($\downarrow$). We report the difference of each method from the baseline performance. Hence, negative differences are better (unlearning works).
@@
PURGE maintains the model utility regardless of model size. To investigate how well our unlearning method scales with model size, we apply Purge to three checkpoints from the Qwen‑2.5 family (0.5B, 1.5B, and 3B parameters) and evaluate them with a subset of the benchmarks introduced in Section 5.1. The results are summarized in Figure 3. Across all three forgetting probes (FB, QA, and AA), we observe a monotonic incline in performance as the parameter count increases. This trend indicates that larger models retain unwanted information more stubbornly and are therefore harder to unlearn, consistent with recent evidence (Carlini et al., 2022) that memorization in LLMs scales superlinearly with size. In contrast, the plot on the right of Figure 3 shows that useful capabilities are largely preserved after PURGE, independent of model scale, which is in line with what we expect to happen based on Theorem 2. All five utility metrics remain stable, indicating that PURGE effectively removes targeted knowledge without harming general downstream performance, even with millions of additional parameters.
6 Discussion
Limitations. Despite its advantages, PURGE has some limitations. Its reward operates at the level of surface token suppression, meaning that although adversarial evaluations demonstrate strong robustness, it remains unclear whether targeted knowledge is fully erased from internal representations or merely rendered difficult to retrieve. This distinction between representational removal and output-level suppression is a broader unresolved challenge in LLM unlearning. In addition, while the synthetic forget corpus substantially reduces token requirements and improves scalability, its effectiveness depends on the quality of entity extraction and probe coverage; missing salient concepts may lead to incomplete forgetting. Although our multi-LLM robustness study (Appendix B.3) shows that PURGE does not depend on a specific proprietary teacher model, corpus construction remains a critical and potentially limiting component.
Reward Design. PURGE employs a minimal binary reward that assigns a value of 1 when no forbidden tokens appear and 0 otherwise, prioritizing deterministic verifiability and eliminating the need for external reward models. When combined with GRPO’s group-relative advantage normalization, this sparse signal is sufficient to produce stable learning dynamics and consistent suppression behavior. However, the simplicity of this design also imposes constraints: it does not directly capture semantic paraphrases beyond extracted entities, nor does it explicitly reason about latent representations, and its sparsity may slow optimization and increase the number of training iterations. Future work could explore richer yet still verifiable reward formulations, such as embedding-level similarity constraints or completion-wide penalties to better approximate semantic forgetting while preserving verifiability, stability and interpretability.
Unlearning Evaluation. Robust evaluation remains one of the most fundamental challenges in LLM unlearning. Existing benchmarks primarily assess observable suppression at inference time rather than the true absence of latent knowledge inside the model’s representations and its output probability distribution. Forget-set accuracy alone cannot guarantee that information has been fully removed, and while adversarial testing provides stronger evidence, exhaustive coverage of all elicitation strategies is infeasible. In theory, retraining a model from scratch without the targeted data is often regarded as the gold standard for validating unlearning. However, for frontier-scale models, this approach is computationally prohibitive, making it impractical for routine evaluation. Consequently, a fundamental gap persists between benchmark-based metrics and the operational guarantees required for real-world LLM unlearning, highlighting the need for more principled and scalable validation frameworks.
Future Directions. While PURGE demonstrates strong empirical performance, several important research directions remain open. First, it is important to apply PURGE to batch unlearning, which more closely reflects realistic deployment scenarios in which multiple data points or concepts must be removed simultaneously. Second, evaluating PURGE in the context of LRMs is a critical next step. Many existing unlearning methods degrade substantially when explicit reasoning or chain-of-thought processes are introduced. Determining whether PURGE maintains its effectiveness under such conditions is essential for assessing its applicability to next-generation models. Finally, integrating external knowledge bases may offer a principled approach to refining the forget set, thereby improving the robustness and accuracy of the entity extraction process.
@@ 7 Conclusion
In this work, we presented PURGE, a novel framework for targeted LLM unlearning that that reframes forgetting as a verifiable optimization problem and leverages GRPO to achieve reliable information removal. Our theoretical analysis provides formal guarantees of information suppression and high-probability bounds showing that forgetting generalizes beyond the prompts seen during training, while protecting overall model quality. Empirical results on the RWKU benchmark show that PURGE outperforms most SoTA methods in fluency and adversarial robustness, while using up to $\times$46 fewer tokens per target compared to SoTA approaches. Overall, our findings highlight the promise of treating unlearning as a verifiable reinforcement learning problem, offering a scalable, efficient, and theoretically grounded direction for reliable LLM unlearning. We believe this paradigm opens new avenues for safe model unlearning, regulatory compliance, and controllable model behavior.
Reproducibility Statement
We provide all resources necessary to reproduce our results. Our implementation of PURGE includes detailed instructions for environment setup and dependency management, as well as executable scripts that enable full reproduction of each experiment from scratch, and will be released upon publication at: https://github.com/strzar/purge. All datasets used in our experiments are publicly available. We report all model hyperparameters in Appendix C.2. For our base models, we use publicly available HuggingFace checkpoints (e.g., Phi-3-Mini-4K-Instruct), and we release our PURGE-unlearned model checkpoints on HuggingFace at: https://huggingface.co/collections/strzara/purge.
Ethics Statement
Our research focuses on improving machine unlearning to support privacy protection, regulatory compliance, and safer deployment of LLMs. All experiments rely exclusively on publicly available datasets, and our benchmarks contain only synthetic or publicly available, non-sensitive information about public figures. While machine unlearning can, in principle, be misused (e.g., to manipulate model behavior), we emphasize its responsible application as a tool for compliance and user protection. Our method does not introduce new capabilities that expand the risk profile of existing LLMs; instead, it aims to improve their responsible, controlled, and compliant use.
Acknowledgements
This work was supported by the Foundation Esslingen and the Munich Center for Machine Learning (MCML). We are grateful to Prof. Dr. Athanassios Skodras and Efstratios Vamvourellis for their valuable feedback and insightful discussions throughout the development of this research. We also thank Felix Steinbauer for his technical support during the project’s initial phases.
@@ A.3 Putting Theorem 1 in GRPO
Notice that this does not mean that GRPO, after unlearning, has a zero probability of emitting forbidden tokens. This means that at the limit, this probability tends to zero, indicating that GRPO is a suitable approach to unlearning.
@@ A.4 Proof of Theorem 2
Since the KL penalty in the GRPO objective explicitly bounds the average per‐prompt KL divergence, we can pull the square‐root bound outside.
@@ A.5 Regret‐to‐Retraining Theorem
Gist of Theorem 3: If you run GRPO for $T$ iterations with a standard $1/\sqrt{t}$ learning‐rate schedule, then on average your task‐loss will be within $O(1/\sqrt{T})$ of the loss achieved by fully retraining from scratch on the retained data, meaning that after a modest number of updates, PURGE’s performance on the kept tasks is essentially as good as a complete retrain, at a tiny fraction of the cost. Notice that this theorem holds; however, we cannot measure it empirically since we are not able to train an LLM from scratch. Therefore, we do not include it in the main content of the paper. We invite researchers to develop empirical demonstrations of this theorem in LLM unlearning and to critically evaluate whether retraining from scratch truly constitutes the gold standard for unlearning in LLMs.
@@
We cast GRPO as mirror descent on the space of token probability distributions, using the KL divergence as the Bregman divergence. Concretely, each update solves a proximal optimization of the form
@@ Appendix B Additional Experiments
B.1 TOFU Benchmark Results
To further evaluate PURGE’s generalization capabilities, we extended our experiments to the TOFU benchmark. Only minor adjustments to the forget–corpus construction were required, confirming that PURGE does not rely on RWKU-specific assumptions and generalizes to alternative unlearning settings. We follow the OpenUnlearning (Dorna et al., 2025) framework and reproduce all baselines on Llama-3.2-1B-Instruct, using the evaluation protocol and reference implementations described in the GitHub repository provided by OpenUnlearning (specifically the results reported in open-unlearning/docs/repro.md).
Table 2 summarizes the results across three key metrics: Forget Quality, Forget Truth Ratio, and Model Utility. Forget Quality and Forget Truth Ratio, closer to 1 indicates stronger forgetting, while a higher Model Utility indicates better preservation of model capabilities. PURGE remains competitive with contemporary approaches (e.g., GradDiff, RMU, UNDIAL), despite operating under a conceptually different mechanism. Although not state-of-the-art in TOFU, PURGE exhibits stable, robust performance, highlighting future opportunities for optimization and deeper theoretical analysis. In summary, these additional results demonstrate that PURGE maintains strong performance on TOFU with minimal methodological adjustments, underscoring its generality and revealing promising avenues for future refinement.
@@ B.1 TOFU Benchmark Results
B.2 Extended Results on Unlearning Finetuning Dynamics
Detailed Reward Analysis
Figure 5 shows the complete reward trajectories for each unlearning-target training run. At each training step, we record the mean reward across all model completions. Initially, most curves begin at a low mean reward, indicating that the model does not yet adhere to our defined reward function (see Equation 13) and then steadily rise, demonstrating that the model progressively learns to satisfy our custom reward criteria and hence unlearns the target.
Figure 5: Training reward trajectories for each unlearning target plotted against the global training step. Each of the 20 subplots shows the full evolution of the model’s ability to track the reward function for a specific target throughout its training run, highlighting differences in convergence speed and stability across experiments. Note that Theorem 1 is satisfied on all targets.
Detailed KL Divergence Analysis
Figure 6: KL divergence trajectories for each unlearning target plotted against the global training step. Each of the 20 subplots shows the evolution of the KL divergence during its respective training run, illustrating variations in divergence-reduction dynamics and convergence behavior across different unlearning targets. Notice that Theorem 2 is satisfied for all targets.
B.3 Robustness to the Choice of NER-Constructing LLM
With these complementary experiments, we evaluated whether our unlearning method maintains stable performance when the forget corpus NER extraction step is performed by different LLMs, including both proprietary and open-source systems. Using the same unlearning target (Stephen King) and the full evaluation suite reported in the main paper, we substituted the NER-constructing model with several state-of-the-art alternatives. Table 3 summarizes the results. Across all metrics, we observe that the unlearning pipeline behaves consistently regardless of which LLM is used for NER extraction. While some models yield marginally stronger performance on isolated metrics, no single model functions as a critical dependency. These results reinforce the robustness of our approach and demonstrate that the method does not rely on any specific proprietary model for success.
Table 3: Robustness of our unlearning pipeline to the choice of NER-constructing LLM. Performance remains stable across proprietary and open-source models, showing no critical dependency on any specific LLM. Arrows denote whether higher ($\uparrow$) or lower ($\downarrow$) values are better. Unlearning target: Stephen King.
Forget Set ↓ Neighbor Set ↑ MIA Set Utility Set ↑
FB QA AA FB QA FM ↑ RM ↓ GA RA FAC TRU FLU
Claude 4.5 Sonnet .464 .250 .497 .700 .603 -2.615 -2.452 .667 .284 .149 .360 6.940
Deepseek R1 .268 .200 .397 .467 .511 -2.619 -2.538 .643 .309 .415 .180 6.696
GPT-5 .339 .288 .558 .567 .623 -2.616 -2.450 .667 .309 .149 .380 6.957
Gemini 3 .464 .338 .489 .667 .596 -2.613 -2.449 .667 .333 .152 .340 6.949
Llama 3.1 405B .339 .271 .558 .567 .646 -2.617 -2.447 .667 .321 .154 .340 6.924
@@ C.1 Baseline Details
User: You are an AI Assistant who is supposed to forget about Stephen King and provide answers without its knowledge, as if you never knew about it. Don’t tell anyone that you unlearned anything.
@@
Building on the rejection templates from the TOFU (Maini et al., 2024) benchmark, we curate 100 negatively labelled prompts and fine‑tune by minimizing
@@ C.2 Hyperparameter Selection
We tuned the number of epochs and the learning rate ($\eta$) per training stage to balance stability (avoiding catastrophic drift from the base Phi-3 model) and sufficient adaptation to each objective. In Table 4 below, we summarize the choices and their rationale.
@@
Overall, we set the epoch count to 3 to simplify scheduling and monitoring, and we scaled $\eta$ inversely with the expected gradient volatility and the risk of catastrophic forgetting at each stage.
@@ D.1 RWKU Dataset Sets
The goal of the forget set is to measure the unlearning efficacy—i.e., whether the model ceases to recall or use the specific knowledge designated for removal. It is partitioned into three splits, which test different aspects of unlearning:
@@
More specifically, the AA split comprises 9 adversarial attack types. We give a few examples:
@@
The MIA set is used to detect whether the model still betrays membership of target fragments in training data, serving as a privacy proxy. It is comprised of the following splits:
@@
The utility set is used to quantify the side effects of unlearning on broader capabilities beyond local neighborhood knowledge. We use well-known benchmarks that test different LLM capabilities: (1) General Ability (Gen): MMLU (5-shot accuracy); (2) Reasoning Ability (Rea): BBH with 27 subtasks, CoT with 3-shot prompts (Exact Match); (3) Truthfulness (Tru): TruthfulQA MC1 (6-shot accuracy); (4) Factuality (Fac): TriviaQA (6-shot F1); and (5) Fluency (Flu): AlpacaEval, weighted bi-/tri-gram entropy.
@@ D.2 Metrics
Measures the longest common subsequence overlap between the model’s output and the reference. Forget Set: Lower is better ($\downarrow$), less overlap means the target knowledge is removed. Neighbor Set: Higher is better ($\uparrow$), we want the model to retain nearby knowledge.
@@ Appendix F PURGE Unlearning Qualitative Examples
To illustrate the practical impact of our PURGE unlearning procedure, we present a series of qualitative examples drawn from three well-known public figures. Each example contrasts the model’s original (“Base Model”) output with its corrected (“PURGE”) output, highlighting how targeted removal of spurious knowledge yields more accurate and robust responses and showcasing the method’s versatility in real-world scenarios.